# Subtype-Aware Registration of Longitudinal Electronic Health Records

### v1

@@ 2 Method for Subtype-Aware Registration of EHR
    Stopping Criterion. We employ stopping criteria to ensure efficient convergence and prevent excessive computation. Based on the initial clustering $\{\mathcal{C}_{1}^{(1)},\ldots,\mathcal{C}_{K^{(1)}}^{(1)}\}$, we evaluate the average Silhouette coefficient $s(K^{(1)})$ using Equation (1). The maximum number of iterations $I$ is set as:
    $I=\left\{\begin{array}[]{ll}I_{\text{small}}&\text{if }s(K^{(1)})>\tau,\\ I_{\text{large}}&\text{if }s(K^{(1)})\leq\tau,\end{array}\right.$
    where $\tau$ is a predefined threshold.
    We terminate the iterations when:
    $\left\{K^{(t)}=K^{(t-1)}\text{ and }t>I\right\}\quad\text{or}\quad\left\{t>2I\right\}.$
    The condition $t>2I$ provides a safeguard to ensure termination even if convergence is not achieved within the expected number of iterations.
2.1 Step 3: Finalization
Upon termination at iteration $t=t_{\text{final}}$, we address the data points not selected during the last iteration. For each cluster $k$, consider the set of unselected points: $\mathcal{C}_{k}^{(t_{\text{final}})}\setminus\mathcal{S}_{k}^{(t_{\text{final}})}$, where $\mathcal{S}_{k}^{(t_{\text{final}})}$ denotes the subset of selected points within cluster $k$ after the final iteration.
For each unselected point $j\in\mathcal{C}_{k}^{(t_{\text{final}})}\setminus\mathcal{S}_{k}^{(t_{\text{final}})}$, find the optimal shift $\delta_{j}^{*}$ and assign it to the most appropriate cluster center by solving:
$\delta_{j}^{*},\ k^{\prime*}=\mathop{\rm arg\min}_{\delta_{j}\in\Lambda,\ k^{\prime}=1,\ldots,K^{(t_{\text{final}})}}\left\|\mathbf{\Omega}_{j,\delta_{j}}-\bar{\boldsymbol{\gamma}}_{k^{\prime}}^{(t_{\text{final}})}\right\|.$
This assigns $j$ to $\mathcal{S}_{k}^{(t_{\text{f}inal})}$ and shift state that best represents it.
The final output of the algorithm is the number of clusters $\widehat{K}^{t_{\rm final}}$, the set of shift states, and clusters for all subjects:
$\{\delta_{i}^{*}:i=1,2,\ldots,N\}\quad\text{and}\quad\{\mathcal{S}_{k}^{(t_{\text{final}})}\}_{k\in\widehat{K}^{(t_{\text{final}})}}.$
@@ 2.1 Step 3: Finalization
1:Longitudinal data $\{(t_{ik},X_{i}(t_{ik}))\mid i=1,\ldots,N;\ k=1,\ldots,T_{i}\}$, Potential shift positions $\Lambda=\{l_{0},l_{1},\ldots,l_{m}\}$, Maximum number of clusters $M$, predefined threshold $\tau$, Parameters $\alpha$, $I_{\text{small}}$, $I_{\text{large}}$, penalization parameter $\lambda$.
2:Optimal shift states $\{\delta_{i}^{*}\mid i=1,2,\ldots,N\}$ and clusters $\{\mathcal{C}_{k}^{(t_{\text{final}})}\}$
3:Introduce internal knots $\{\tau_{1},\ldots,\tau_{q}\}$ in $[T_{\min},T_{\max}]$
@@
$\widehat{\boldsymbol{\beta}}_{i,l}=\mathop{\rm arg\min}_{\boldsymbol{\beta}_{i,l}}\left\{\sum_{\begin{subarray}{c}k=1\\ T_{\min}\leq t_{ik}-l\leq T_{\max}\end{subarray}}^{T_{i}}\left(X_{i}(t_{ik}-l)-\sum_{j=1}^{p}\beta_{ij}B_{j,3}(t_{ik})\right)^{2}+\lambda\sum_{j=1}^{p}\beta_{ij}^{2}\right\}$
@@
1:Initialize time shifts $\delta_{i}^{(0)}=0$ for all $i$. Set iteration counter $t\leftarrow 1$
@@
4:         Perform clustering with $k$ clusters on $\{\mathbf{\Omega}_{i,\delta_{i}^{(t-1)}}\}_{i=1}^{N}$
5:         Compute average Silhouette coefficient $s(k)$ using (1)
@@
7:     Select optimal number of clusters: $K^{(t)}=\mathop{\rm arg\max}_{k}s(k)$; obtain clusters $\{\mathcal{C}_{1}^{(t)},\ldots,\mathcal{C}_{K^{(t)}}^{(t)}\}$
8:     for each cluster $k=1$ to $K^{(t)}$ do
9:         Compute centroid: $\bar{\boldsymbol{\beta}}_{k}^{(t)}=\sum_{i\in\mathcal{C}_{k}^{(t)}}\mathbf{\Omega}_{i,\delta_{i}^{(t-1)}}/|\mathcal{C}_{k}^{(t)}|$
10:         Calculate distances $d_{i}=\left\|\mathbf{\Omega}_{i,\delta_{i}^{(t-1)}}-\bar{\boldsymbol{\beta}}_{k}^{(t)}\right\|$ for $i\in\mathcal{C}_{k}^{(t)}$
11:         Select subset $\mathcal{S}_{k}^{(t)}$ of points with smallest $\left\lceil\alpha|\mathcal{C}_{k}^{(t)}|\right\rceil$ distances
12:         Recompute centroid: $\bar{\boldsymbol{\gamma}}_{k}^{(t)}=\sum_{i\in\mathcal{S}_{k}^{(t)}}\mathbf{\Omega}_{i,\delta_{i}^{(t-1)}}/\frac{1}{|\mathcal{S}_{k}^{(t)}|}$
@@
14:     for each subject $i=1$ to $N$ do
15:         Update time shift: $\delta_{i}^{(t)}=\mathop{\rm arg\min}_{\delta\in\Lambda}\left\|\mathbf{\Omega}_{i,\delta}-\bar{\boldsymbol{\gamma}}_{k}^{(t)}\right\|^{2}$, where $k$ is the cluster containing subject $i$
@@
17:     Update $t\leftarrow t+1$
18:     Check Stopping Criteria
19:     if $K^{(t)}=K^{(t-1)}$ and $t>I$ then
20:         Terminate Iteration
21:     else if $t>2I$ then
22:         Terminate Iteration (Hard Stop)
23:     end if
24:until Termination condition is met
Algorithm 3 Step 3: Iterative Clustering and Time Shift Adjustment
@@
2:     for each unselected point $j\in\mathcal{C}_{k}^{(t_{\text{final}})}\setminus\mathcal{S}_{k}^{(t_{\text{final}})}$ do
@@
$\delta_{j}^{*},\ k^{\prime*}=\mathop{\rm arg\min}_{\delta_{j}\in\Lambda,\ k^{\prime}=1,\ldots,K^{(t_{\text{final}})}}\left\|\mathbf{\Omega}_{j,\delta_{j}}-\bar{\boldsymbol{\gamma}}_{k^{\prime}}^{(t_{\text{final}})}\right\|.$
Then reassign $j$ to cluster $\mathcal{S}_{k^{\prime}*}^{(t_{\text{final}})}$.

### v2

@@ 3 Method for Subtype-Aware Registration of EHR
    Stopping Criterion.
    We employ stopping criteria to ensure efficient convergence and to prevent excessive computation. Given an initial clustering $\{\mathcal{C}_{1}^{(1)},\ldots,\mathcal{C}_{K^{(1)}}^{(1)}\}$, we evaluate the average silhouette coefficient $s(K^{(1)})$ using the equation presented earlier in the section. When $s(K^{(1)})>\tau$ for some predefined threshold $\tau$, we set the maximum number of iterations to be 1 and stop the iterations; otherwise, we run more iterations as follows.
    At each iteration $h\geq 2$, let $K^{(h)}$ denote the number of clusters selected by maximizing the average silhouette coefficient over a candidate range, and let $s_{[2]}^{(h)}=\max_{\begin{subarray}{c}k\in\{2,\dots,M\}\\ k\neq K^{(h)}\end{subarray}}s^{(h)}(k)$ denote the second-highest silhouette score among all $k\neq K^{(h)}$. The iterative process terminates when the following combined condition is satisfied:
    $\left\{K^{(h)}=K^{(h-1)}\;\text{ and }\;s^{(h)}(K^{(h)})\leq s^{(h-1)}(K^{(h-1)})\;\text{ and }\;s_{[2]}^{(h)}\leq s_{[2]}^{(h-1)}\right\}\text{ or }\left\{h>Num_{stop}\right\},$
    where $Num_{stop}$ is a hard cap on the number of iterations.
    This unified condition ensures the iteration breaks if the clustering solution stabilizes, if the initial quality is already sufficient, or if a maximum number of iterations is reached.
Step 3: Finalization
Upon termination at iteration $h=h_{\text{final}}$, we address the data points not selected during the last iteration. For each cluster $k$, consider the set of unselected points: $\mathcal{C}_{k}^{(h_{\text{final}})}\setminus\mathcal{S}_{k}^{(h_{\text{final}})}$, where $\mathcal{S}_{k}^{(h_{\text{final}})}$ denotes the subset of selected points within cluster $k$ after the final iteration.
For each unselected point $j\in\mathcal{C}_{k}^{(h_{\text{final}})}\setminus\mathcal{S}_{k}^{(h_{\text{final}})}$, find the optimal shift $\delta_{j}^{*}$ and assign it to the most appropriate cluster center by solving:
$\delta_{j}^{*},\ k^{\prime*}=\mathop{\rm arg\min}_{\delta_{j}\in\Lambda,\ k^{\prime}=1,\ldots,K^{(h_{\text{final}})}}\left\|\mathbf{\Omega}_{j,\delta_{j}}-\bar{\boldsymbol{\gamma}}_{k^{\prime}}^{(h_{\text{final}})}\right\|.$
This assigns $j$ to $\mathcal{S}_{{k^{\prime}}^{*}}^{(h_{\text{f}inal})}$ and shift state that best represents it.
The final output of the algorithm is the number of clusters $\widehat{K}^{h_{\rm final}}$, the set of shift states, and clusters for all subjects:
$\{\delta_{i}^{*}:i=1,2,\ldots,N\}\quad\text{and}\quad\{\mathcal{S}_{k}^{(h_{\text{final}})}\}.$
@@
1:Longitudinal data $\{(t_{ir},X_{i}(t_{ir}))\mid i=1,\ldots,N;\ r=1,\ldots,T_{i}\}$, Potential shift positions $\Lambda=\{l_{0},l_{1},\ldots,l_{m}\}$, Maximum number of clusters $M$, predefined threshold $\tau$, Parameters $\alpha$, penalization parameter $\lambda$.
2:Optimal shift states $\{\delta_{i}^{*}\mid i=1,2,\ldots,N\}$ and clusters $\{\mathcal{C}_{k}^{(h_{\text{final}})}\}$
3:Introduce knots $\{\kappa_{1},\ldots,\kappa_{q+8}\}$ in $T_{\min},T_{\max}]$
@@
$\widehat{\boldsymbol{\beta}}_{i,l}=\mathop{\rm arg\min}_{\boldsymbol{\beta}_{i,l}}\left\{\sum_{\begin{subarray}{c}r=1\\ T_{\min}\leq t_{ir}+l\leq T_{\max}\end{subarray}}^{T_{i}}\left(X_{i}(t_{ir})-\beta_{i,l,0}-\sum_{j=2}^{p}\beta_{i,l,j}B_{j,3}(t_{ir}+l)\right)^{2}+\lambda\sum_{j=2}^{p}\beta_{i,l,j}^{2}\right\}$
@@
1:Initialize time shifts $\delta_{i}^{(0)}=0$ for all $i$. Set iteration counter $h\leftarrow 1$
@@
4:   Perform clustering with $k$ clusters on $\{\mathbf{\Omega}_{i,\delta_{i}^{(h-1)}}\}_{i=1}^{N}$
5:   Compute the average silhouette coefficient $s(k)$
@@
7:  Select the optimal number of clusters: $K^{(h)}=\mathop{\rm arg\max}_{k}s(k)$; obtain clusters $\{\mathcal{C}_{1}^{(h)},\ldots,\mathcal{C}_{K^{(h)}}^{(h)}\}$
8:  for each cluster $k=1$ to $K^{(h)}$ do
9:   Compute centroid: $\bar{\boldsymbol{\beta}}_{k}^{(h)}=\sum_{i\in\mathcal{C}_{k}^{(h)}}\mathbf{\Omega}_{i,\delta_{i}^{(h-1)}}/|\mathcal{C}_{k}^{(h)}|$
10:   Calculate distances $d_{i}=\left\|\mathbf{\Omega}_{i,\delta_{i}^{(h-1)}}-\bar{\boldsymbol{\beta}}_{k}^{(h)}\right\|$ for $i\in\mathcal{C}_{k}^{(h)}$
11:   Select subset $\mathcal{S}_{k}^{(h)}$ of points with smallest $\left\lceil\alpha|\mathcal{C}_{k}^{(h)}|\right\rceil$ distances
12:   Recompute centroid: $\bar{\boldsymbol{\gamma}}_{k}^{(h)}=\sum_{i\in\mathcal{S}_{k}^{(h)}}\mathbf{\Omega}_{i,\delta_{i}^{(h-1)}}/|\mathcal{S}_{k}^{(h)}|$
@@
14:  for each subject $i=1$ to NN do
15:   Update time shift: $\delta_{i}^{(h)}=\mathop{\rm arg\min}_{\delta\in\Lambda}\left\|\mathbf{\Omega}_{i,\delta}-\bar{\boldsymbol{\gamma}}_{k}^{(h)}\right\|^{2}$, where $k$ is the cluster containing subject $i$
@@
17:  Check Stopping Criteria:
18:  if $h=1$ and $s(K^{(1)})>\tau$ then
19:   Terminate
20:  end if
21:  $K^{(h)}\leftarrow\arg\max_{k}s^{(h)}(k)$; $s^{(h)}\leftarrow s^{(h)}(K^{(h)})$; $s_{[2]}^{(h)}\leftarrow\max_{\begin{subarray}{c}k\in\{2,\dots,M\}\\ k\neq K^{(h)}\end{subarray}}s^{(h)}(k)$
22:  if $K^{(h)}=K^{(h-1)}$, $s^{(h)}(K^{(h)})\leq s^{(h-1)}(K^{(h-1)})$ and $s_{[2]}^{(h)}\leq s_{[2]}^{(h-1)}$} or $h>Num_{stop}$} then
23:   Terminate
24:  end if
25:until Termination condition is met
Algorithm 3 Step 3: Assign Unselected Points and Finalize
@@
2:  for each unselected point $j\in\mathcal{C}_{k}^{(h_{\text{final}})}\setminus\mathcal{S}_{k}^{(h_{\text{final}})}$ do
@@
$\delta_{j}^{*},\ k^{\prime*}=\mathop{\rm arg\min}_{\delta_{j}\in\Lambda,\ k^{\prime}=1,\ldots,K^{(h_{\text{final}})}}\left\|\mathbf{\Omega}_{j,\delta_{j}}-\bar{\boldsymbol{\gamma}}_{k^{\prime}}^{(h_{\text{final}})}\right\|.$
Then reassign $j$ to cluster $\mathcal{S}_{{k^{\prime}}^{*}}^{(h_{\text{final}})}$.